OpenStudy (anonymous):

solve a2 + 8a + 12 = 0 the solution(s) is/are a = __? or there is no solution

OpenStudy (saifoo.khan):

a = -2 a = -6

OpenStudy (anonymous):

\[a^2+8a+12=0\]\[(a+6)[(a+2)=0\] \[(a+6)=0\]\[a=-6\]\[(a+2)=0\]\[a=-2\]

OpenStudy (anonymous):

thanks california.mathematics for the arrival to the answer

OpenStudy (anonymous):

uwc

OpenStudy (anonymous):

so the anser is -6, -2 right

OpenStudy (anonymous):

plug it in and check which one solution

OpenStudy (anonymous):

i don't know exactly how to do that

OpenStudy (anonymous):

\[-6^2+8(-6)+12=36-48+12=0\] \[-2^2+8(-2)+12=4-16+12=0\] yes:: -6 and -2 is solution